evise a procedure to dilute part of the 0.200 mol/L solution to make 20 mL of 0.150 mol/L solution. (hint:  use the dilution formula).  Write a sentence that describes how much of your initial 0.200 mol/L solution needs to be used, and how much water needs to be added to make this solution



1
Expert's answer
2021-01-25T04:29:28-0500

When the solution is diluted, the number of the moles of the solute remains constant:

"n_1 = n_2" .

The number of the moles of the solute in a solution is the product of its molar concentration and the volume of the solution:

"c_1V_1 = c_2V_2" .

This equation is called a dilution formula. In this formula, the concentration 1 is 0.200 mol/L, the concentration 2 is 0.150 mol/L and the volume 2 is 20 mL. From this data, we can find the volume 1:

"V_1 = \frac{c_2V_2}{c_1}= \frac{0.150·20}{0.200} = 15" mL.

Therefore, in order to prepare the 20 mL of 0.150 mol/L solution, we need 15 mL of 0.200 mol/L solution. Finally, the volume of water one needs to add is 20-15=5 mL.


Answer: 15 mL of the initial 0.200 mol/L solution and 5 mL of water are needed to prepare 20 mL of 0.150 mol/L solution.